# Audio–Visual Speech Enhancement: Architectural Design and Deployment Strategies

### v2

@@ Abstract
This paper presents a real-time wireless 5G cloud-based audio-visual speech enhancement (AVSE) framework that integrates convolutional neural networks (CNNs) for acoustic feature extraction and OpenCV-based facial embedding for visual processing. The extracted multimodal features are fused through a long short-term memory (LSTM) network to reconstruct clean speech from noisy input. Experimental evaluations across multiple network configurations—Wi-Fi, 4G, 5G, Ethernet, and AWS cloud—demonstrate that only 5G and direct Ethernet links consistently satisfy the coherence condition ($t_{\text{comm}}\leq t_{\text{chunk}}$) for continuous streaming. By applying an 80× compression factor, the payload per 40 ms audio-video chunk is reduced from approximately 0.3 MB to 70 KB with minimal quality loss, enabling reliable transmission under bandwidth constraints. Results further reveal a trade-off between algorithmic latency and enhancement performance: smaller models and shorter input windows lower processing time but reduce speech clarity. The proposed architecture achieves end-to-end latency of approximately 5 s with strong noise reduction and speech intelligibility, demonstrating the feasibility of deploying multimodal speech enhancement in cloud-assisted, low-latency wireless environments.
@@ I Introduction
In many real-world acoustic environments, speech signals are severely degraded by additive background noise, reverberation, and competing speakers, which substantially reduce intelligibility and perceived quality. Conventional audio-only speech enhancement techniques aim to recover a clean speech signal ($x[n]$) from a noisy observation ($y[n]=x[n]+d[n]$), where ($d[n]$) denotes noise or interference. However, when the signal-to-noise ratio (SNR) is low or overlapping talkers are present, these audio-only approaches often struggle to reliably recover intelligible speech.
Human listeners naturally exploit not only auditory cues but also visual information — such as lip movements and facial articulations — to compensate for challenging acoustic conditions. This perceptual phenomenon is well known (e.g., the McGurk effect) and underscores the value of multimodal input in speech perception. Visual speech cues remain unaffected by acoustic masking and therefore offer complementary information that is particularly valuable when the acoustic channel is corrupted.
Audio-visual speech enhancement (AVSE) builds on this insight by jointly leveraging both audio and visual modalities (for example, microphone audio plus camera capture of the speaker’s lips) to enhance or recover target speech. Deep-learning based AVSE systems typically extract acoustic features (e.g., log-mel spectrograms, STFT magnitudes) and visual features (e.g., lip region embeddings via CNNs), and then fuse them in a neural network to generate enhanced speech. Surveys (e.g., Michelsanti et al. 2021 [michelsanti2021overview]) have shown that multimodal AVSE systems often outperform audio-only baselines, especially in low-SNR or multi-speaker conditions.
@@
Despite these algorithmic and modelling advances, deploying AVSE in real-world applications introduces a distinct set of system-level challenges. Wearable devices such as hearing aids or smart glasses must operate under tight constraints of latency, power consumption, and wireless connectivity. They also need to handle streaming audio and video data in real time. Offloading heavy deep‐learning inference to a remote cloud or edge server emerges as a feasible solution when the local device is resource‐constrained.
In this context, the advent of next-generation wireless networks (e.g., 5G NR) with high bandwidth and ultra-low latency offers a promising platform for real-time AVSE deployment. For example, recent demonstrations of 5G/IoT enabled AVSE for multimodal hearing aids illustrate the viability of high-data-rate, low-latency uplinks combined with cloud-based enhancement [gupta20235g]. By harnessing the 5G link and off-loading computation to the cloud, devices can remain light, efficient and power-conserving, while still benefiting from powerful multimodal models.
However, the full stack of implementation trade‐offs—including uplink/downlink latency, wireless reliability under mobility, compute placement (device vs edge vs cloud), streaming bandwidth for visual + audio, and real-time intelligibility/quality outcomes—remains under-explored in AVSE research. The present work addresses this gap by proposing and implementing a wireless 5G-based cloud architecture for AVSE. Our system captures lip-region video and noisy speech on a wearable front‐end, transmits the data via a 5G uplink to a remote cloud server, performs multimodal deep-learning AVSE processing, and streams back the enhanced speech to the device. We evaluate the system in terms of end-to-end latency, wireless throughput, model inference time, and speech intelligibility/quality under realistic noisy and multi-speaker conditions. In so doing, we extend the state of the art from model-centric AVSE research to a full-stack, deployment-oriented system suitable for mobile and wearable real-time use.
The remainder of this paper is structured as follows. Section II reviews related work on AVSE, wireless/5G off‐load implementations and deployment strategies. Section III presents the system design and implementation of our 5G cloud-based AVSE pipeline. Section IV reports experimental evaluation results and discusses findings and practical implications. Finally, Section V concludes the paper and outlines future research directions.
@@ II Related Work
Research in audio-visual speech enhancement (AVSE) has progressed substantially over the past decade, particularly with the proliferation of deep learning and multimodal approaches. Early work in audio-visual speech separation, such as the seminal model by Ephrat et al. (“Looking to Listen at the Cocktail Party”), demonstrated that a joint audio‐visual network significantly outperforms audio‐only baselines in multi‐speaker and noisy conditions by leveraging lip‐motion cues [ephrat2018looking]. Subsequently, surveys such as Michelsanti et al. (2021) [michelsanti2021overview] provided a systematic overview of the field, highlighting key components such as acoustic feature extraction, visual feature encoding, fusion strategies (early vs intermediate vs late), and standard objective functions for AVSE/AVSS.
In more algorithm‐centric developments, attention has turned to how to fuse modalities effectively and design architectures that can operate in real‐time or near‐real‐time. For example, Xu and Hao [xu2021multi] proposed the attentional multi‐layer fusion convolution network (AMFFCN) to apply soft‐threshold attention at each layer of the decoder to select informative modalities. More recently, Gao and Grauman introduced the VisualVoice framework, which learns cross-modal embeddings (face appearance plus audio) for improved speech separation in unconstrained video scenes [gao2021visualvoice].
In terms of system implementation and deployment, particularly for wearable or real‐world applications, researchers have begun to tackle not just the algorithmic enhancement but also the architectural trade-offs of where the processing is done (device, edge, or cloud). Adeel et al. (2018) [adeel2018real] envisioned a 5G IoT enabled lip-reading driven hearing‐aid framework that off-loads AV data to the cloud, employing lightweight encryption and aiming for round‐trip latency targets on the order of a few milliseconds. On the dataset and baseline front, Michelsanti’s survey [michelsanti2021overview] underscores the importance of dataset design and evaluation metrics (e.g., STOI, PESQ, SI-SDR) when comparing AVSE systems.
Despite algorithmic advances, few works systematically compare deployment strategies (cloud vs edge vs device) under realistic network and computing constraints. To fill this gap, our work develops a comprehensive AVSE architecture (CNN + LSTM fusion) and performs a comparative analysis across deployment scenarios under Ethernet, Wi-Fi, 4G and 5G conditions. Although, prior work has established the foundational value of adding visual cues to speech enhancement; explored a wide variety of fusion architectures and benchmarks; and begun to address deployment challenges, The full spectrum of practical trade‐offs—end‐to‐end latency, network delay, compute location, and real‐world system performance—remains less explored, motivating the present implementation and experimental‐study contribution.
III Proposed Solution
@@
Figure 1: A block diagram of enabling real-world COG-MHEAR service on terminal devices.
@@
Figure 2 describes the latency of the processing of the COG-MHEAR service. As shown in the system model, the client continuously collects voice and video information and sends it to the server. Assume that the length of the audio-video chunk sent by the client each time is $t_{\text{chunk}}$, that is, the client will send out audio and video after collecting a chunk of $t_{\text{chunk}}$. After receiving the data, the server will first perform preprocessing, such as feature extraction of video frames, and then store the processed data in the buffer. Since our algorithm tracks lips to help audio enhancement, the input of the algorithm must have a specific length $t_{i}$ to ensure sufficient visual features. When there are more than $t_{i}$ data in the buffer, the server will start executing the AVSE algorithm. Then, the server processes the incoming audio and video data sequentially, and we must set a sufficient buffer interval $t_{\Delta}$ to ensure that it can complete each process, which is given as
@@
where $t_{a}$ is the algorithm processing latency. The enhanced audio is stored in the output buffer, and the server begins feeding the enhanced audio back to the client when there is enough data in the output buffer. It should be noted that since $t_{i}$ is usually much larger than $t_{\Delta}$, if we want the initial audio to be enhanced by this algorithm, there will be an inevitable $t_{i}$-length waiting time for data, causing a high delay. So we play a trick, that is, we do not process the audio at the beginning of $t_{i}-t_{\Delta}$, thereby reducing the overall delay to
@@
IV Experimental Results
In this experiment, the auditory information undergoes a transformation via convolutional neural networks (CNNs), which can distill the audio patterns. The visual information is processed through the OpenCV library, which is an open-source computer vision and machine learning software library, to extract facial features from the face embeddings. It is implemented using the pre-trained Haar feature-based cascade classifiers, which are underpinned also by CNNs. Subsequently, the algorithm employs a fusion strategy where the extracted auditory and facial features coalesce, serving as inputs to a sophisticated long shortterm memory (LSTM) network. This step is crucial for integrating temporal dynamics and ensuring a coherent feature representation. The convergence of these modalities culminates in a series of FC layers, which tunes the composite features to reconstruct the target outputs. This framework provides a foundational blueprint, while the explicit architectural configuration (i.e., layers, neurons, etc.) can vary depending on the requirements and data pre-processing.
@@ IV Experimental Results
We measured the latency of a variety of networks in real-world environments. We record audio at a sampling rate of $16$ kHz and captured video frames at a resolution of $380\times 640$ pixels at a frame rate of $25$ fps. We set $t_{\text{chunk}}=40$ ms, congruent with the temporal extent of a single video frame. This configuration resulted in a raw data payload of approximately $0.3$ MB per audio-video chunk. Figure 3 shows the latency profiles observed across a suite of $100$ round-trip data transmissions. In our experimental setup, a desktop-based server and a laptop-based client were utilized to simulate a realistic user scenario. The communication between server and client was facilitated through Wi-Fi (802.11n, i.e., Wi-Fi 4), 4G, 5G, and a direct Ethernet cable link, respectively. We also tried the Amazon Web Services, which hosts the server codes, and the laptop-based client accesses it through the same Wi-Fi. The results suggest that only 5G network and direct Ethernet cable link could support a coherent service experience, i.e., satisfying $t_{\text{comm}}\leq t_{\text{chunk}}$ where $t_{\text{chunk}}=40$ ms in our experiments. Note that 4G and 5G network capacities may differ in different regions and with different carriers while choosing different AWS locations may also result in different latencies. We used Wi-Fi from the King’s Buildings, The University of Edinburgh (Edinburgh, UK), and for 4G and 5G network testing used a private network of Cambridge Wireless (Cambridge, UK). Using new Wi-Fi protocols, such as 802.11ac (Wi-Fi 5) or 802.11ax (Wi-Fi 6), can dramatically reduce the latency. In areas with poor signal coverage, we can compress the raw audio-video data to reduce the transmission load, obtaining reliable and consistent quality of service. Figure 4 shows the data size versus compression factors. It can be seen that proper compression of data can greatly reduce the data size and the required data rate, especially in the range of 80-100 compression factors. At a compression factor of 80, the data can be reduced to $\tilde{7}0$ KB with a trivial effect on the quality of the video frames. We also used a compression factor of 80 in our real-world demonstration in Cambridge to ensure smooth service, and experiments proved that the impact of other speech enhancement properties could be ignored and we could still achieve amazing noise reduction.
@@
IV-A Algorithm Processing Latency
The communication latency has little effect on user experience delay ($t_{\text{delay}}$) but is critical to the coherence of feedback audio. And the real-world measurements show the ability to meet the demand using either the latest 5G or WI-Fi 6. In order to reduce the user experience latency, the most critical thing is to reduce the processing time of the algorithm, which can be achieved by reducing the input chunk size and the model size. However, since our algorithm needs to track lip vibrations in order to extract the target speaker’s voice from the mixed noise, decreasing the size of the input chunk usually incurs a performance loss. Figure 5 shows the algorithm processing latency versus the input chunk size in terms of frames, where one frame is $40$ ms. It can be seen that the processing delay is almost proportional to the input size. Theoretically, if we process each frame individually, then the processing latency will be less

### v4

# Audio-Visual Speech Enhancement: Architectural Design and Deployment Strategies

@@ Abstract
Real-time audio-visual speech enhancement (AVSE) is a key enabler for immersive and interactive multimedia services, yet its performance is tightly constrained by network latency, uplink capacity, and computational delay. This paper presents the design, deployment, and evaluation of a complete cloud–edge–assisted AVSE system operating over a public 5G edge network. The system integrates CNN-based acoustic enhancement and OpenCV-based facial feature extraction with an LSTM fusion network to preserve temporal coherence, and is deployed on a Vodafone-compatible AWS Wavelength edge cloud. Through extensive stress testing, we analyze end-to-end performance under varying network load and adaptive multimedia profiles. Results show that compute placement at the network edge is critical for meeting real-time coherence constraints, and that uplink capacity is often the dominant bottleneck for interactive AVSE services. Only 5G and wired Ethernet consistently satisfied the required communication delay bound for uncompressed audio-video chunks, while aggressive compression reduced payload sizes by up to 80× with negligible perceptual degradation, enabling robust operation under constrained conditions. We further demonstrate a fundamental trade-off between processing latency and enhancement quality, where reduced model complexity lowers delay but degrades reconstruction performance in low-SNR scenarios. Our findings indicate that public 5G edge environments can sustain real-time, interactive AVSE workloads when network and compute resources are carefully orchestrated, although performance margins remain tighter than in dedicated infrastructures. The architectural insights derived from this study provide practical guidelines for the design of delay-sensitive multimedia and perceptual enhancement services on emerging 5G edge-cloud platforms.
@@ I Introduction
Speech signals captured in real-world environments are often severely degraded by background noise, reverberation, and competing speakers, leading to reduced intelligibility and perceived quality. Traditional audio-only speech enhancement techniques struggle to recover clean speech in low signal-to-noise ratio (SNR) settings or with overlapping talkers, because audio cues alone provide limited information under such conditions. In contrast, humans naturally exploit visual cues—such as lip movements and facial articulations—to disambiguate speech in noise, demonstrating that visual information is robust to acoustic masking and can significantly improve speech understanding in adverse environments. Audio-visual speech enhancement (AVSE) systems leverage this multimodal insight by fusing acoustic and visual features through deep neural networks, often outperforming audio-only approaches in low-SNR or multi-speaker scenarios.
@@
Despite these advances in multimodal modeling, deploying AVSE in real-time applications introduces a distinct set of system-level challenges. Wearable and mobile devices—such as hearing aids, smart glasses, or assistive communication tools—must operate under strict latency, power, and connectivity constraints while processing continuous streams of synchronized audio and video. Offloading compute-intensive deep-learning inference to remote servers emerges as a viable solution for resource-limited devices, but this exposes the system to network latency, jitter, and variable throughput that can degrade performance or violate temporal coherence requirements.
The advent of 5G and edge cloud computing offers a promising platform for addressing these challenges. Multi-access edge computing (MEC) brings cloud capabilities closer to end users by deploying compute resources at or near the radio access network (RAN), thereby reducing data transport delay and backbone congestion. Standards bodies such as ETSI define MEC as an environment where applications can run at the edge of 4G/5G networks with ultra-low latency and high bandwidth [4]. Edge cloud platforms like AWS Wavelength embed cloud infrastructure within telecommunications networks, enabling low-latency access to compute and storage services directly from 5G devices [2]. While prior work has demonstrated the theoretical potential of MEC for latency-sensitive services, real-world empirical studies of audio-visual multimedia workloads over public 5G edge deployments remain limited, especially at the intersection of multimodal processing, network variability, and end-to-end performance.
Audio-visual multimedia workloads are inherently demanding: they combine temporally synchronized audio and video streams, require sustained throughput for high-resolution visual data, and must maintain tight end-to-end latency to preserve perceptual coherence. Even modest delay variation or packet loss can disrupt synchronization, degrade quality, and undermine user experience in interactive settings. These challenges are further exacerbated in mobile environments, where wireless channel conditions and user mobility introduce variability that centralized cloud architectures cannot easily accommodate.
In this work, we address these gaps by presenting a full-stack, deployment-oriented study of audio-visual multimedia delivery and processing over a public 5G edge cloud, using AVSE as a representative real-time workload. We design and implement an AVSE pipeline that captures synchronized audio and lip-region video on a wearable front end, transmits data via a 5G uplink to an edge cloud server, performs multimodal deep learning-based enhancement, and streams enhanced speech back to the device. The system is deployed on a commercial 5G edge platform, and its performance is evaluated in terms of end-to-end latency, throughput, packet loss, compute utilization, and enhancement quality under realistic network conditions and adaptive stress profiles.
Our contributions are as follows:
  • We present a real-world deployment and end-to-end evaluation of audio-visual multimedia processing over a public 5G edge cloud, extending AVSE research from model-centric prototypes to practical systems.
  • We introduce stress-adaptive multimedia profiles that systematically explore throughput, latency, and quality trade-offs across heterogeneous network conditions.
  • We quantitatively measure network and system performance—including uplink throughput, round-trip latency, resource utilization, and packet loss—to identify constraints on real-time operation in public 5G environments.
  • We derive practical insights and design guidelines for latency-sensitive multimedia services on 5G edge clouds, highlighting the importance of edge proximity, adaptive profiles, and multi-modal synchronization for real-time QoE.
The remainder of this paper is organized as follows. Section II reviews related work on audio-visual speech enhancement, edge/cloud deployment strategies, and 5G-enabled multimedia systems. Section III presents the proposed audio-visual speech enhancement architecture and its deployment over a public 5G edge cloud, including system design and latency modeling. Section IV reports experimental results from real-world deployments over Ethernet, Wi-Fi, 4G, and public 5G networks, and analyzes throughput, latency, reliability, and compute utilization under stress-adaptive multimedia profiles. Section V discusses algorithmic processing latency and model optimization for real-time operation. Finally, Section VI concludes the paper and outlines directions for future research.
@@ II Related Work
Research on audio–visual speech enhancement (AVSE) has advanced significantly with the adoption of deep learning and multimodal fusion architectures. Early seminal work demonstrated that combining audio with visual cues such as lip motion significantly improves speech separation and enhancement under noisy and multi-speaker conditions compared to audio-only methods. For example, Ephrat et al. jointly model audio and visual modalities for speaker separation, establishing a foundation for multimodal speech processing [3]. Subsequent models such as multimodal deep convolutional networks (e.g., AVDCNN) validate the effectiveness of integrated audio and visual feature extraction and joint decoding for enhanced speech quality [9].
Beyond model design, more recent AVSE studies have explored enhanced fusion strategies (e.g., dense connectivity and gated fusion) and real-time capable architectures, while also considering robust dataset construction and domain-specific tasks such as spatial audio processing or feature alignment [13].
While the majority of AVSE research focuses on improving model performance in controlled conditions, fewer studies tackle the challenges of real-time deployment where network transport, latency constraints, and computational placement are critical. In [8] the authors showcased a proof-of-concept implementation of the world’s first 5G and Internet of Things (IoT) enabled multi-modal hearing aid (MM HA) prototype. This integrates an innovative 5G cloudradio access network (C-RAN) and IoT based transceiver model for real-time audio-visual speech enhancement (AVSE). Our work evaluates AVSE performance under realistic network conditions (Ethernet, Wi-Fi, 4G, and 5G), highlighting trade-offs between communication delay, model complexity, and enhancement effectiveness.
Another related direction investigates 5G-enabled IoT and wearable systems, where audiovisual data are streamed to cloud servers for enhancement and secure processing, addressing both latency and privacy concerns in emerging assistive applications [1]. Beyond AVSE, broader work on mobile edge computing (MEC) and 5G performance analysis provides insights into system design for interactive multimedia services, where computation offloading and resource allocation strategies are critical to support low-latency processing required by live audiovisual applications [10]. Additionally, work on wireless offloading and wearable edge computing suggests that high-throughput, low-latency links can effectively support the real-time transport of high-resolution visual streams for vision-driven workloads such as object detection, further emphasizing the potential of advanced wireless architectures for AVSE systems [12].
III Problem Formulation: Audio–Visual Speech Enhancement
Audio–visual speech enhancement (AVSE) aims to recover a clean target speech signal by jointly exploiting acoustic and visual observations of the speaker. Let the discrete-time noisy audio signal captured at the microphone be
$y[n]=x[n]+d[n],$ (1)
where $x[n]$ denotes the clean target speech and $d[n]$ represents additive noise and interference. In adverse acoustic environments, recovering $x[n]$ from $y[n]$ alone is ill-posed, particularly at low signal-to-noise ratio (SNR) or in the presence of competing speakers. Visual cues, however, remain largely invariant to acoustic corruption and therefore provide complementary information for speech recovery.
Let $\mathbf{a}_{t}\in\mathbb{R}^{F_{a}}$ denote the audio feature vector extracted from $y[n]$ over time frame $t$ (e.g., log-mel or STFT magnitude features), and let $\mathbf{v}_{t}\in\mathbb{R}^{F_{v}}$ denote the corresponding visual feature vector extracted from the speaker’s facial region (e.g., lip motion embeddings). The AVSE problem is formulated as learning a nonlinear mapping
$\hat{\mathbf{x}}_{t}=f_{\theta}(\mathbf{a}_{1:t},\mathbf{v}_{1:t}),$ (2)
where $f_{\theta}(\cdot)$ denotes a deep neural network with parameters $\theta$, and $\hat{\mathbf{x}}_{t}$ is the estimated clean speech representation at time $t$. The model is trained by minimizing the reconstruction loss
$\min_{\theta}\;\mathcal{L}=\sum_{t=1}^{T}\left\|\hat{\mathbf{x}}_{t}-\mathbf{x}_{t}\right\|_{2}^{2},$ (3)
where $\mathbf{x}_{t}$ denotes the clean target speech features and $T$ is the sequence length.
IV Proposed Solution
IV-A Model Architecture and Feature Fusion
The proposed AVSE architecture consists of parallel audio and visual processing streams followed by feature fusion and temporal modeling, as illustrated in Fig. 1.
Figure 1: Audio-Visual Speech Enhancement AI model architecture.
The audio stream employs a convolutional neural network (CNN) encoder
$\mathbf{z}_{t}^{(a)}=\mathcal{E}_{a}(\mathbf{a}_{t}),$ (4)
while the visual stream extracts facial features using a vision front-end based on OpenCV and CNN feature extractors,
$\mathbf{z}_{t}^{(v)}=\mathcal{E}_{v}(\mathbf{v}_{t}).$ (5)
The modality-specific embeddings are fused via concatenation,
$\mathbf{z}_{t}=\left[\mathbf{z}_{t}^{(a)}\;\|\;\mathbf{z}_{t}^{(v)}\right],$ (6)
where $\cdot\|\cdot$] denotes vector concatenation. Temporal dependencies are modeled using a long short-term memory (LSTM) network,
$\mathbf{h}_{t}=\mathrm{LSTM}(\mathbf{z}_{t},\mathbf{h}_{t-1}),$ (7)
which ensures coherent speech reconstruction across time. A stack of fully connected (FC) layers decodes the enhanced speech features,
$\hat{\mathbf{x}}_{t}=\mathcal{D}(\mathbf{h}_{t}).$ (8)
IV-B Implementation
@@
Figure 2: A block diagram of enabling real-world AVSE service on terminal devices.
@@
Figure 3 describes the latency of the processing of the AVSE service. As shown in the system model, the client continuously collects voice and video information and sends it to the server. Assume that the length of the audio-video chunk sent by the client each time is $t_{\text{chunk}}$, that is, the client will send out audio and video after collecting a chunk of $t_{\text{chunk}}$. After receiving the data, the server will first perform preprocessing, such as feature extraction of video frames, and then store the processed data in the buffer. Since our algorithm tracks lips to help audio enhancement, the input of the algorithm must have a specific length $t_{i}$ to ensure sufficient visual features. When there are more than $t_{i}$ data in the buffer, the server will start executing the AVSE algorithm. Then, the server processes the incoming audio and video data sequentially, and we must set a sufficient buffer interval $t_{\Delta}$ to ensure that it can complete each process, which is given as
@@
where $t_{a}$ is the algorithm processing latency. The enhanced audio is stored in the output buffer, and the server begins feeding the enhanced audio back to the client when there is enough data in the output buffer. It should be noted that since $t_{i}$ is usually much larger than $t_{\Delta}$, if we want the initial audio to be enhanced by this algorithm, there will be an inevitable $t_{i}$-length waiting time for data, causing a high delay. In order to mitigate the significant delays we do not process the audio at the beginning of $t_{i}-t_{\Delta}$, thereby reducing the overall delay to
@@
Figure 4: End-to-end system architecture for 5G edge cloud multimedia delivery.
V Experimental Results
In this experiment, the auditory information undergoes a transformation via convolutional neural networks (CNNs), which can distill the audio patterns. The visual information is processed through the OpenCV library, which is an open-source computer vision and machine learning software library, to extract facial features from the face embeddings. It is implemented using the pre-trained Haar feature-based cascade classifiers, which are underpinned also by CNNs. Subsequently, the algorithm employs a fusion strategy where the extracted auditory and facial features coalesce, serving as inputs to a sophisticated Long- Short Term Memory (LSTM) network. This step is crucial for integrating temporal dynamics and ensuring a coherent feature representation. The convergence of these modalities culminates in a series of FC layers, which tunes the composite features to reconstruct the target outputs. This framework provides a foundational blueprint, while the explicit architectural configuration (i.e., layers, neurons, etc.) can vary depending on the requirements and data pre-processing.
@@ V Experimental Results
We measure the latency of a variety of networks in real-world environments. We record audio at a sampling rate of $16$ kHz and captured video frames at a resolution of $380\times 640$ pixels at a frame rate of $25$ fps. We set $t_{\text{chunk}}=40$ ms, congruent with the temporal extent of a single video frame. This configuration yields a raw data payload of approximately $0.3$ MB per audio-video chunk. Figure 5 shows the latency profiles observed across a suite of $100$ round-trip data transmissions. In our experimental setup, a desktop-based server and a laptop-based client were utilized to simulate a realistic user scenario. The communication between server and client was facilitated through Wi-Fi (802.11n, i.e., Wi-Fi 4), 4G, 5G, and a direct Ethernet cable link, respectively. We also tried the Amazon Web Services over internet as well as over Vodafone 5G public network whose architecture is depicted in Figure 4. The results suggest that only private 5G network and direct Ethernet cable link could support a coherent service experience, i.e., satisfying $t_{\text{comm}}\leq t_{\text{chunk}}$ where $t_{\text{chunk}}=40$ ms in our experiments. Note that 4G and 5G network capacities may differ in different regions and with different carriers while choosing different AWS locations may also result in different latencies. We used Wi-Fi from the King’s Buildings, The University of Edinburgh (Edinburgh, UK), and for 4G and 5G network testing used a private network of Cambridge Wireless (Cambridge, UK) and Vodafone 5G public network. Using new Wi-Fi protocols, such as 802.11ac (Wi-Fi 5) or 802.11ax (Wi-Fi 6), can dramatically reduce the latency.
In areas with poor signal coverage, we can compress the raw audio-video data to reduce the transmission load, obtaining reliable and consistent quality of service. Figure 6 shows the data size versus compression factors. It can be seen that proper compression of data can greatly reduce the data size and the required data rate, especially in the range of 80-100 compression factors. At a compression factor of 80, the data can be reduced to $\approx 70$ KB with an insignificant effect on the quality of the video frames. We also used a compression factor of 80 in our real-world demonstration in Cambridge to ensure smooth service, and experiments proved that the impact of other speech enhancement properties could be ignored and we could still achieve considerable noise reduction.
(a) Wi-Fi (802.11n)
(b) 4G
(c) 5G
(d) Ethernet cable
(e) Amazon Web Services
(f) AWS Wavelength over Vodafone 5G Cellular Network
@@
The next section describes extensive Vodafone 5G public network stress testing with different data compression profiles.
V-A Vodafone 5G network
This section describes the experimental environment used to evaluate real-time audio-visual multimedia delivery over a public 5G network integrated with edge cloud infrastructure. We first outline the network characteristics and edge cloud configuration, followed by details of the user equipment setup, stress profile design, and performance metrics used for analysis.
The experiments were conducted over a commercial Vodafone 5G network operating in the United Kingdom’s mid-band spectrum (3.4–3.6 GHz), which provides a balance of capacity and coverage suitable for real-world edge deployments. Targeted network characteristics include an end-to-end latency of less than 15 ms for edge-hosted applications and sustained upstream bandwidth in excess of 50 Mbps. Typical downstream 5G throughput is expected to range between 100–200 Mbps, significantly higher than legacy 4G under ideal conditions, thus accommodating bandwidth-intensive audio-visual workloads.
The edge cloud environment was hosted using AWS Wavelength Zones, which embed compute and storage resources directly within the operator’s network to minimize transport distance. In this deployment, application servers were placed within a private subnet accessible only through the operator’s carrier gateway. A bastion host provided controlled secure shell access for development and system management, ensuring that compute resources remained isolated from direct public Internet exposure. General-purpose instance types supported by Wavelength include t3.medium, t3.xlarge, and r5.2xlarge, along with GPU-accelerated options such as g4dn.2xlarge; this work uses the t3.medium instance type to represent a typical baseline for edge compute.
User equipment consisted of a Google Pixel 8 smartphone used as the primary 5G access device. To facilitate logging and control from a standard development environment, the phone’s mobile data connection was shared with a laptop via Wi-Fi tethering, effectively creating a local hotspot. All multimedia payloads were generated by client applications on the tethered laptop and transmitted over the 5G access network to the edge application servers. This setup ensured realistic traffic behavior on the mobile access link while providing flexibility for measurement and control.
To systematically explore system performance under varying network loads and processing demands, we defined a set of AVSE profiles. These profiles vary in video frame rate, spatial resolution, JPEG compression quality, and audio frame configuration, reflecting a range from very low to extreme bandwidth usage. Table I summarizes the profile parameters used in this study.
TABLE I: AVSE Multimedia Profile Configurations
Profile Category Video FPS Resolution JPEG Quality Audio Frame (ms) Bandwidth Class
[…]
5A Audio-only 20 Very low
5B Audio-only 10 Low
Performance was assessed using several key metrics. Round-trip time (RTT) was measured between the user equipment and the edge application server, excluding application processing overhead, to characterize pure network latency. Throughput measurements captured sustained uplink and downlink bandwidth during multimedia transmission. Packet loss was tracked at the network level to quantify reliability, and jitter was analyzed to understand timing variation affecting real-time media. Finally, resource utilization on the edge compute node, particularly host CPU usage, was monitored to evaluate processing overhead under different profile loads. Together, these metrics provide a comprehensive view of network, transport, and system performance supporting real-time multimedia over 5G edge cloud deployments.
The analysis below focuses on throughput scalability, latency behavior, system resource utilization, and reliability across AVSE profiles ranging from audio-only baselines to extreme audio-visual configurations.
(a) Average downlink throughput.
(b) Average RTT (network latency only).
(c) Average host CPU utilization.
(d) Average packet loss.
Figure 7: AWS edge-cloud (Wavelength) over Vodafone 5G network performance metrics across different AVSE profiles under increasing load.
Fig. 7(a) presents the average downlink throughput for the five stress classes. Throughput increases monotonically from the audio-only (AO) baseline to the AV Stress profile, demonstrating that the proposed profile design provides controlled and predictable scaling of network load. Low and medium stress profiles (AV Low and AV Medium) increase throughput gradually, enabling graceful adaptation under constrained network conditions. In contrast, the AV High and AV Stress profiles introduce a sharp increase in bandwidth demand, with the latter exceeding 13 Mbps on average, representing the upper operational limit of the pipeline. These results validate the ability of the system to span a wide range of deployment scenarios using a single unified profile design.
Fig. 7(b) shows the corresponding network latency measured as average RTT, excluding AVSE processing overhead. RTT remains relatively stable for AO, AV Low, and AV Medium profiles (approximately 54–59 ms), indicating that moderate increases in video bitrate do not significantly impact network delay when edge cloud deployment is used. However, latency increases for AV High and peaks for AV Stress at approximately 73 ms, suggesting the onset of queuing and scheduling effects under sustained high throughput. While still acceptable for real-time audio-visual processing, these results highlight the reduced latency margin at extreme configurations and the importance of adaptive profile selection.
Fig. 7(c) illustrates the host CPU utilization across stress levels. Audio-only processing requires minimal resources (approximately 5% CPU usage), while AV Low and AV Medium profiles operate between 10–13%, representing a practical region for sustained deployment. The AV High profile increases CPU usage to approximately 18%, and the AV Stress profile exceeds 21%, reflecting the combined cost of high-rate video encoding and network transmission. These results indicate that computational load scales predictably with stress level and that mid-range profiles provide the best efficiency–performance trade-off.
Fig. 7(d) reports the average packet loss observed across stress profiles. Packet loss remains low for AO through AV High profiles, demonstrating the robustness of the pipeline under typical operating conditions. The AV Stress profile exhibits higher loss and variance, confirming that it is primarily suitable for stress testing rather than continuous real-time operation. Notably, AV Medium and AV High profiles maintain low packet loss while delivering significantly higher throughput, reinforcing their suitability for practical deployments.
Overall, the results demonstrate that while extreme profiles expose system and network limits, mid-range configurations achieve a favorable balance between throughput, latency, resource consumption, and reliability. These findings confirm that edge cloud proximity, combined with adaptive profile selection, is a key enabler for real-time audio-visual speech enhancement in public 5G environments.
TABLE II: Latency report of different AVSE deployments
Network Latency (ms) Notes
Ethernet (Direct) $\sim$5–15 Stable, lowest latency
5G private network (Cambridge Wireless) $\sim$20–35 Satisfies real-time requirements
5G public network (Vodafone + AWS Wavelength, London) 54.4–73 Acceptable for real-time A/V with reduced margin
Wi-Fi 4 (802.11n) $\sim$50–100 Too high for smooth streaming; Wi-Fi 5/6 recommended
4G (Cambridge Wireless) $\sim$60–150 Exceeds latency threshold, inconsistent
AWS over Internet (WebRTC) $\sim$80–200 Region and ISP dependent; cloud adds delay
The latency measurements highlight the strong dependence of real-time audio-visual speech enhancement (AVSE) performance on network architecture rather than access technology alone. Ethernet (direct) provides the lowest and most stable latency ($\approx$5–15 ms), representing the ideal baseline for real-time AVSE and confirming that processing, rather than networking, becomes the dominant factor in wired deployments.
The private 5G network achieves consistently low latency ($\approx$20–35 ms), comfortably satisfying real-time requirements and demonstrating the advantage of localized cores and controlled radio environments for edge-based AVSE services.
Vodafone’s public 5G network with AWS Wavelength in London exhibits moderate latency ($\approx$54–73 ms), which remains acceptable for real-time audio-visual applications, though with reduced margin. This increase reflects the added complexity of public network traversal and edge-cloud integration, yet still significantly outperforms traditional internet-based cloud deployment.
In contrast, Wi-Fi 4 and 4G show higher and more variable latency ($\approx$50–150 ms), often exceeding thresholds for smooth interactive streaming. These results indicate that legacy access technologies and shared radio conditions introduce instability that limits AVSE usability.
Finally, deploying AVSE over the public internet to AWS regions using WebRTC results in the highest and most unpredictable latency ($\approx$80–200 ms), driven by geographic distance, ISP routing, and core network hops. This clearly demonstrates the importance of edge computing for latency-sensitive audio-visual applications.
Overall, the results confirm that edge proximity is the dominant factor enabling real-time AVSE, and that public 5G with edge cloud provides a practical middle ground between private networks and best-effort internet cloud deployments.
V-B Algorithm Processing Latency